### v3

# IIT Gandhinagar Gujarat, India Arxiv

Abstract
Neural radiance fields (NeRF) have revolutionized photorealistic rendering of novel views for 3D scenes. Despite their growing popularity and efficiency as 3D resources, NeRFs face scalability challenges due to the need for separate models per scene and the cumulative increase in training time for multiple scenes. The potential for incrementally encoding multiple 3D scenes into a single NeRF model remains largely unexplored. To address this, we introduce Continual-Neural Graphics Primitives (C-NGP), a novel continual learning framework that integrates multiple scenes incrementally into a single neural radiance field. Using a generative replay approach, C-NGP adapts to new scenes without requiring access to old data. We demonstrate that C-NGP can accommodate multiple scenes without increasing the parameter count, producing high-quality novel-view renderings on synthetic and real datasets. Notably, C-NGP models all $8$ scenes from the Real-LLFF dataset together, with only a $2.2\%$ drop in PSNR compared to vanilla NeRF, which models each scene independently. Further, C-NGP allows multiple style edits in the same network.
Figure 1: (Left) The proposed continual neural graphics primitives (C-NGP) framework to model multiple scenes. Every new scene is assigned a unique pseudo-label and trained sequentially while preserving the previously learned information via generative replay. Any specific scene can be rendered by conditioning C-NGP over the associated pseudo-label. (Right) Our modification to the Instant-NGP architecture [Müller et al.(2022)Müller, Evans, Schied, and Keller] enables multi-scene modeling using pseudo labels.
1 Introduction
Neural Radiance Fields (NeRF) synthesize photorealistic novel views by modeling 3D scenes using Multi-Layer Perceptrons (MLPs) and differentiable volumetric rendering. While powerful, NeRF and its variants typically require per-scene optimization, leading to high training and storage costs. Recent methods address speed [Wang et al.(2021a)Wang, Liu, Liu, Theobalt, Komura, and Wang, Fridovich-Keil et al.(2022)Fridovich-Keil, Yu, Tancik, Chen, Recht, and Kanazawa, Müller et al.(2022)Müller, Evans, Schied, and Keller], inference efficiency [Chen et al.(2023c)Chen, Funkhouser, Hedman, and Tagliasacchi, Wan et al.(2023)Wan, Richardt, Božič, Li, Rengarajan, Nam, Xiang, Li, Zhu, Ranjan, et al.], compression [Gordon et al.(2023)Gordon, Chng, MacDonald, and Lucey, Chen et al.(2022)Chen, Xu, Geiger, Yu, and Su, Tang et al.(2022)Tang, Chen, Wang, and Zeng, Shin and Park(2024), Fang et al.(2023)Fang, Xu, Wang, Yang, Wang, and Zhou], scene editing [Wang et al.(2023b)Wang, Wang, Qu, and Qi, Bao et al.(2023)Bao, Zhang, Yang, Fan, Yang, Bao, Zhang, and Cui], and generalization to unseen scenes [He et al.(2023)He, Liang, Xiao, Chen, and Chen, Yu et al.(2021)Yu, Ye, Tancik, and Kanazawa, Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser, Zhang et al.(2022)Zhang, Bi, Sunkavalli, Su, and Xu, Yang et al.(2023a)Yang, Hong, Li, Hu, Li, Lee, and Wang, Chen et al.(2023a)Chen, Yi, Ma, Jia, and Lu]. However, learning multiple distinct scenes within a single NeRF model remains underexplored, with SCARF [Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi] being a recent attempt.
Training separate NeRFs for each scene is impractical for large-scale 3D collections due to resource demands. We propose a continual learning approach that incrementally encodes multiple scenes within a single NeRF while maintaining a fixed parameter count. Our framework learns new scenes sequentially without forgetting previous ones and without having to store past training data. Effective multi-scene modeling requires: (a) photorealistic rendering of all scenes, (b) integration of new scenes without access to old data, (c) prevention of catastrophic forgetting, and (d) constant model size. Moreover, continual learning in NeRFs is uniquely difficult. Unlike image classification, where an image seldom has different labels, NeRF requires scene-specific mappings since the same position and view direction across different scenes produce different outputs.
We introduce C-NGP - Continual Neural Graphics Primitives that leverage fast training and efficient learning of Instant-NGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] by integrating scene conditioning into its hashing mechanism using pseudo-labels to accommodate multiple scenes into NGP’s parameter space. This simple yet effective conditioning enables multi-scene optimization, allowing the model to learn new scenes while preserving high-fidelity rendering of previously encountered scenes. C-NGP uses generative-replay to accommodate new scenes without requiring access to training data from previously learned scenes. Thus, C-NGP is practical for real-world use, allowing multiple 3D scenes to be stored and shared in a single network checkpoint, where the new scenes can be added incrementally without creating separate models. Interestingly, even vanilla NeRF can be adapted to the proposed setting, but with far higher computational overhead (12 days of training time vs. 8 hours for C-NGP) and lower performance.
Our main contributions are: 1) C-NGP, the first fixed-size NeRF to continually encode multiple scenes, achieving only a $2.2\%$ PSNR drop on Real-LLFF [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar] compared to scene-specific vanilla NeRF; 2) training without access to previous data via generative replay; 3) an efficient scene-conditioning mechanism in Instant-NGP’s hash grid; 4) support for multi-view consistent style editing across scenes; 5) competitive performance over continual learning baselines [Chung et al.(2022)Chung, Lee, Baik, and Lee, Cai and Müller(2023), Po et al.(2023)Po, Dong, Bergman, and Wetzstein] and concurrent methods [Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi] in quality, training efficiency, and adaptability to new scenes.
2 Related Work
Single-Scene Neural Radiance Fields. NeRF encodes scene geometry and appearance but requires long per-scene optimization. Alternatives like Plenoxels [Fridovich-Keil et al.(2022)Fridovich-Keil, Yu, Tancik, Chen, Recht, and Kanazawa] and Instant-NGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] improve efficiency using spherical harmonics and multi-resolution hash encoding. Other extensions enable faster inference [Chen et al.(2023c)Chen, Funkhouser, Hedman, and Tagliasacchi, Wan et al.(2023)Wan, Richardt, Božič, Li, Rengarajan, Nam, Xiang, Li, Zhu, Ranjan, et al.], scene editing [Wang et al.(2023b)Wang, Wang, Qu, and Qi, Bao et al.(2023)Bao, Zhang, Yang, Fan, Yang, Bao, Zhang, and Cui], or rendering from sparse images [Lee et al.(2024)Lee, Choi, Kim, Kim, and Cho, Yang et al.(2023b)Yang, Pavone, and Wang, Wang et al.(2023a)Wang, Chen, Loy, and Liu]. Another line of work aims to generalize NeRF to unseen scenes without per-scene training [Yu et al.(2021)Yu, Ye, Tancik, and Kanazawa, Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser, Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su, He et al.(2023)He, Liang, Xiao, Chen, and Chen, Zhang et al.(2022)Zhang, Bi, Sunkavalli, Su, and Xu, Yang et al.(2023a)Yang, Hong, Li, Hu, Li, Lee, and Wang]. These methods use convolutional feature extractors, 3D cost volumes, or attention modules [Park et al.(2019)Park, Florence, Straub, Newcombe, and Lovegrove, Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman] to learn scene priors, enabling inference on new scenes. However, they often require fine-tuning for optimal quality and are not designed for incremental learning, suffering from catastrophic forgetting when adapting to multiple scenes in a shared parameter space. Unlike these approaches, our method focuses on continual learning to incrementally encode distinct scenes without forgetting, targeting compact 3D asset storage rather than generalization. Recently, 3D Gaussian Splatting (3DGS) methods [Kerbl et al.(2023)Kerbl, Kopanas, Leimkuehler, and Drettakis, Chen et al.(2024)Chen, Xu, Zheng, Zhuang, Pollefeys, Geiger, Cham, and Cai] have achieved real-time rendering for single scenes but do not support incremental multi-scene learning. However, these approaches continue to focus on single-scene optimization.
Continual Learning in NeRF. Continual learning enables models to learn new tasks sequentially without forgetting prior knowledge [Robins(1995)]. Techniques include parameter isolation [Mallya and Lazebnik(2018)], regularization [Kirkpatrick et al.(2017)Kirkpatrick, Pascanu, Rabinowitz, Veness, Desjardins, Rusu, Milan, Quan, Ramalho, Grabska-Barwinska, et al., Li and Hoiem(2017)], and replay [Shin et al.(2017)Shin, Lee, Kim, and Kim, Rannen et al.(2017)Rannen, Aljundi, Blaschko, and Tuytelaars]. In NeRF, applying continual learning is challenging due to scene-specific representations, as the same 3D point and view direction yield different densities and colors across scenes. Most existing methods focus on single-scene scenarios, addressing appearance or geometry changes over time [Chung et al.(2022)Chung, Lee, Baik, and Lee, Cai and Müller(2023), Po et al.(2023)Po, Dong, Bergman, and Wetzstein, Zhang et al.(2023a)Zhang, Li, Chen, and Xu]. For example, MEIL-NeRF [Chung et al.(2022)Chung, Lee, Baik, and Lee] and CLNeRF [Cai and Müller(2023)] reconstruct scenes from partial scans but are limited to single-scene forgetting. SLAM-based methods [Sucar et al.(2021)Sucar, Liu, Ortiz, and Davison, Chen et al.(2023b)Chen, Chen, Wang, Zhang, Guo, Shan, and Wang, Deng et al.(2024)Deng, Shen, Qin, Wang, Zhao, Wang, Wang, and Chen] assume past data availability for keyframe optimization, unsuitable for multi-scene settings without stored data.
Multi-Scene Continual Learning in NeRF. Encoding multiple distinct scenes in a single NeRF model is largely unexplored. A concurrent work, SCARF [Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi], uses a hyper-network to generate scene-specific weights, but this increases the parameter count with each new scene. In contrast, our C-NGP method leverages generative replay [Shin et al.(2017)Shin, Lee, Kim, and Kim] to learn new scenes without storing past data, maintaining a fixed parameter count. By integrating scene conditioning into Instant-NGP’s [Müller et al.(2022)Müller, Evans, Schied, and Keller] hashing mechanism, C-NGP achieves high-fidelity rendering and fast convergence, outperforming continual learning baselines [Chung et al.(2022)Chung, Lee, Baik, and Lee, Cai and Müller(2023)] and competitive performance with SCARF [Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi] in multi-scene scenarios.
3 Method
We present C-NGP, a framework to incrementally model multiple scenes with the same set of parameters that is traditionally used for modeling a single scene in an NGP network. We begin with a brief overview of learning neural radiance fields and the multi-resolution hash encoding (MHE) technique. Next, we explore the conditional capabilities of Instant-NGP and describe how generative replay [Shin et al.(2017)Shin, Lee, Kim, and Kim] can be integrated into conditional Instant-NGP. This extension enables incremental learning of new scenes while preserving information from previously encountered scenes.
3.1 Background
Neural Radiance Fields. Mildenhall et al. [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] introduced NeRF, a method for synthesizing novel views from a set of input images using a multi-layer perceptron (MLP) that maps a 5D input—3D spatial coordinates $(x,y,z)$ and viewing direction $(\theta,\phi)$—to the density $\sigma$ and color $c$ of a point. Points are sampled along camera rays $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$, where $\mathbf{o}$ is the camera origin, $\mathbf{d}$ is the ray direction, and $t$ is the sampled depth. Rendering is performed via volumetric integration [Max(1995)], computing the color of a ray as $\hat{C}(r)=\sum_{i=1}^{N}w_{i}c_{i}$, where the weights $w_{i}=T_{i}\alpha_{i}$ depend on transmittance $T_{i}$ and opacity $\alpha_{i}=(1-e^{-\sigma_{i}\delta_{i}})$, with $\delta_{i}$ being the distance between adjacent samples. The model is trained by minimizing the squared error between predicted and ground-truth pixel colors: $\mathcal{L}=\sum_{r\in\mathcal{R}}||\hat{C}(r)-C(r)||_{2}^{2}$, where $\mathcal{R}$ denotes the set of rays. Inputs to the MLP are encoded using sinusoidal positional encodings to better capture high-frequency scene details.
Multi-Resolution Hash Encoding. Müller et al. [Müller et al.(2022)Müller, Evans, Schied, and Keller] introduced the concept of multi-resolution hash encoding (MHE) to enhance both the reconstruction accuracy and training efficiency of neural radiance fields (NeRFs) while maintaining minimal computational overhead. Unlike the traditional positional encoding used in NeRF, MHE employs a trainable multi-level 3D grid structure. The position of a sampled point is encoded by interpolating the features $(\mathbb{R}^{F})$ located at the vertices of the grid cell containing the point. These interpolated features are then fed into an MLP network, predicting both the density $(\sigma)$ and the point’s color $(c)$.
The grid is organized into $L$ levels, each containing a hash table of size $T\times F$, where $T$ represents the number of features and $F$ is the dimensionality of each feature. To efficiently map the grid coordinates, a spatial hash function $h(x)$, based on the approach of [Teschner et al.(2003)Teschner, Heidelberger, Müller, Pomerantes, and Gross], is utilized:
$h(x)$ $=\left(\bigoplus_{i=1}^{d}x_{i}\pi_{i}\right)\text{ mod }T$ (1)
Here, $\oplus$ denotes the bitwise XOR operation, and $\pi_{i}$ are distinct large prime numbers.
3.2 Continual Neural Graphics Primitive (C-NGP)
This section introduces the C-NGP framework that incorporates scene-conditioning and continual learning in the Instant-NGP network [Müller et al.(2022)Müller, Evans, Schied, and Keller].
Conditioning: To condition Instant-NGP, we use pseudo integer labels $C\in$1, 2, 3, …} to differentiate scenes. Figure 1 illustrates the information flow. To get the Neural Hash Encoding (NHE) of coordinates, we combine the scene coordinates ($x$, $y$, $z$) with the pseudo label ($C$). Additionally, the pseudo labels undergo sinusoidal positional encoding (PE) [Tancik et al.(2020)Tancik, Srinivasan, Mildenhall, Fridovich-Keil, Raghavan, Singhal, Ramamoorthi, Barron, and Ng], denoted as $\psi\in\mathbb{R}^{2\alpha}$, where $\alpha=2$ represents the number of frequency components used. The neural hash encoding (NHE) and encoded pseudo labels $(\psi)$ are concatenated and passed through a single-layer MLP to obtain density ($\sigma$) and scene features. The predicted scene feature is then used to compute the color for each input coordinate ($x$, $y$, $z$). Specifically, the scene feature is concatenated with the encoded viewing direction ($SHE$) and encoded pseudo label ($\psi$), forming the input to a two-layer MLP that predicts the final color values. The viewing direction is encoded using spherical harmonics encoding (SHE) [Ramamoorthi and Hanrahan(2001)], further refining the rendering process. We visualize how neural hashing helps segregate scenes within a shared representation space upon scene-conditioning in the supplementary material.
Continual learning: While conditioning helps non-conflicting scene representations, continual learning accommodates new, unseen scenes incrementally. Therefore, the proposed C-NGP involves training Instant-NGP (with scene conditioning) in a continual learning paradigm as follows. Each scene is encoded into conditional Instant-NGP using a pseudo-label. Before encoding a new scene, we first render all previously observed scenes using the camera parameters of the new scene. These rendered images are then added to the training set of the new scene, and the model parameters are updated accordingly. This strategy eliminates reliance on previous training datasets while ensuring high-fidelity modeling of all observed scenes. Rendering previously seen scenes with updated camera parameters is particularly effective for datasets like NeRF Synthetic $360^{\circ}$. For other scene types, we use the stored camera parameters for rendering.
3.3 Loss and Regularization
Floaters represent a significant challenge in radiance fields, typically manifesting as disconnected, dense spatial regions near the camera plane. Therefore, apart from the conventional rendering loss $\mathcal{L}_{mse}$, we include two additional regularizations to ensure stable C-NGP training.
Distortion. Following the work of [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman, Wynn and Turmukhambetov(2023)], we used distortion loss for compact point distribution on the camera ray:
$\mathcal{L}_{dist}=\frac{1}{d(r)}\Bigg{(}\sum_{i}w_{i}w_{j}\Bigg{|}\frac{t_{i}+t_{i+1}}{2}-\frac{t_{j}+t_{j+1}}{2}\Bigg{|}+\frac{1}{3}\sum_{i=1}^{N}w_{i}^{2}(t_{i+1}-t_{i})\Bigg{)}$ (2)
where, $d(r)=\frac{\sum_{i=1}^{N}w_{i}t_{i}}{\sum_{i=1}^{N}w_{i}}$ is depth along each ray. The first part of $\mathcal{L}_{dist}$ minimizes the weighted distance between all pairs of interval midpoints. The second part focuses on minimizing the weighted size of each interval. Jointly, the weights on the ray are encouraged to be compact by pulling distance intervals closer by consolidating each weight and minimizing the width of each interval [Barron et al.(2022)Barron, Mildenhall, Verbin, Srinivasan, and Hedman].
Ray entropy. The entropy regularization calculates the entropy of the ray’s distribution [Kim et al.(2022)Kim, Seo, and Han, Bonotto et al.(2024)Bonotto, Sarrocco, Evangelista, Imperoli, and Pretto] for each ray passing through the scene. This involves assessing the uncertainty or randomness in the predicted densities along the ray and avoiding the floaters in the rendered scene. The entropy regularization is described as per Equation 3
$\mathcal{L}_{ent}$ $=\left(-\sum_{i=1}^{N}p(r_{i})log(p(r_{i})\right)$ (3)
Here, $N$ is the number of sampled points on the ray $r$, and $p_{i}$ is the opacity of each sampled point.
Complete loss. To train the C-NGP method, we used the combined loss formulation, as described in Equation 4.
$\mathcal{L}_{total}$ $=\mathcal{L}_{mse}+\lambda_{ent}\mathcal{L}_{ent}+\lambda_{dist}\mathcal{L}_{dist}$ (4)
Here, $\lambda_{ent}$ and $\lambda_{dist}$ are the weights for the regularization. We keep these parameters for training the complete network as $\lambda_{ent}=1e-3$ and $\lambda_{dist}=1e-2$.
4 Experiments
In this section, we perform an exhaustive set of experiments to demonstrate the efficacy of the proposed framework. We start by discussing the datasets, evaluation metrics, and comparison baselines.
Figure 2: Qualitative Comparison on the Nerf Synthetic $360^{\circ}$ dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]. C-NGP is compared with MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] and IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] (both with per-scene fine-tuning), and vanilla NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] with per-scene optimization. C-NGP is observed to model scenes with minimal artifacts compared to those present in the other methods. C-NGP (FT) is fine-tuned for each test scene for five epochs after continual learning.
4.1 Datasets and Evaluation Metrics
We evaluate our model on widely used datasets: NeRF Synthetic $360^{\circ}$ [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng], Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun], and Forward-Facing LLFF [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar]. NeRF Synthetic $360^{\circ}$ consists of eight diverse scenes, each with $100$ training views and $200$ test views at a resolution of $800\times 800$. The images are rendered from either an upper hemisphere or a full hemisphere, and the camera parameters are not necessarily consistent across scenes. Forward-Facing LLFF contains eight real-world scenes captured with a handheld camera in a forward-facing manner. The number of images per scene varies from $20$ to $62$, with a resolution of $1008\times 756$. Tanks and Temples feature five large-scale scenes with complex geometries and real-world objects, each captured at a resolution of $1920\times 1080$. We use the masked version of this dataset for training and testing, following [Liu et al.(2020)Liu, Gu, Zaw Lin, Chua, and Theobalt]. Additionally, we generate a dataset of $22$ scenes using BlenderNeRF [Raafat()]. Each scene is created from freely available 3D object meshes and rendered with $100$ training views and $100$ test views from an upper hemisphere. This dataset is designed to stress test the continual learning setup of C-NGP and analyze the representative upper bound of network parameters. To ensure fair evaluation in a continual learning setup, all scenes in this dataset are rendered with the same camera parameters.
Metrics. We quantitatively evaluate the model performance Peak-Signal-to-Noise-Ratio (PSNR) [psn(1987)], Structural Similarity Index (SSIM) [Wang et al.(2004)Wang, Bovik, Sheikh, and Simoncelli], and Perceptual Score (LPIPS) [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang].
4.2 Training Details
For training C-NGP, we used a batch size of $10,000$ rays and an initial learning rate of $2\times 10^{3}$. The model was trained for $30$ epochs across synthetic and real datasets. C-NGP was conditioned on pseudo labels assigned to each scene to ensure scene-specific representations. The hash table size for Instant-NGP was set to $T=2^{19}$, with a feature size of $F=4$. For the NeRF Synthetic $360^{\circ}$ dataset, we applied the generative replay method [Shin et al.(2017)Shin, Lee, Kim, and Kim], where previously observed scenes were re-rendered using the camera parameters of the new scene. For real-world datasets like LLFF [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar] and Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun], which lack full $360^{\circ}$ coverage, we only store the camera parameters of each scene. This allowed us to generate novel views while avoiding artifacts such as floaters or ghosting, less problematic issues in synthetic datasets with complete $360^{\circ}$ coverage.
Figure 3: Qualitative Comparison on the Real Forward Facing (LLFF) dataset [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar]. We compare C-NGP with MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] and IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] (both with per-scene fine-tuning), and vanilla NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] with per-scene. C-NGP (FT) is fine-tuned for each test scene for five epochs after continual learning.
Method Tanks and Temples
PSNR $\uparrow$ SSIM $\uparrow$
EWC [Kirkpatrick et al.(2017)Kirkpatrick, Pascanu, Rabinowitz, Veness, Desjardins, Rusu, Milan, Quan, Ramalho, Grabska-Barwinska, et al.] + NeRF 15.64 0.420
PackNet [Mallya and Lazebnik(2018)] + NeRF 16.71 0.547
MEIL-NeRF [Chung et al.(2022)Chung, Lee, Baik, and Lee] 17.98 0.580
CLNeRF [Shin et al.(2017)Shin, Lee, Kim, and Kim] 21.30 0.640
SCARF [Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi] 26.78 0.89
C-NGP (ours) 26.93 0.87
Method NeRF Synthetic 360 Real Forward-Facing
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
SCARF [Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi] 30.94 0.94 - 26.44 0.80 -
C-NGP (ours) 29.40 0.94 0.09 22.12 0.66 0.37
Offline Sampling Online Sampling
Datasets PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeRF Synthetic 360 25.211 0.86 0.151 29.40 0.93 0.09
Blender Synthetic 37.735 0.98 0.022 37.734 0.98 0.022
Table 1: (Left) Quantitative comparison of C-NGP on the Tanks and Temple dataset [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] with baselines that combine either conventional continual learning methods with NeRF or introduce continual learning for single scene optimization along with a concurrent work. (Right-Top) Quantitative comparison of C-NGP on NeRF Synthetic [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] and Real LLFF [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar] with concurrent work. (Right-Bottom) Analyzing the quantitative performance on training with (online) and without (offline) generative replay while learning C-NGP. The enhanced performance with online sampling is due to the same camera parameters across all the scenes.
4.3 Experimental Evaluation
Choice of Baselines. Since no existing work incrementally models multiple scenes in a single radiance field, we compare our approach with the most closely related baselines that align, at least partially, with our experimental setup [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su, Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser, He et al.(2023)He, Liang, Xiao, Chen, and Chen, Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi]. We have shown the quantitative analysis results in the supplementary. Among them, SCARF [Wang et al.(2024)Wang, Wang, Wang, Duan, Bao, and Qi] is the closest and most concurrent to our work. It handles multiple scenes, however, by increasing the number of parameters through a global parameter generator, unlike us. We compare the quantitative performance with SCARF based on their reported results in the paper across different datasets, and omit the qualitative comparison due to the unavailability of their code.
Quantitative. We compare the proposed paradigm with continual learning frameworks like Elastic Weight Consolidation (EWC) and PacketNet [Mallya and Lazebnik(2018)] combined with NeRF, and single-scene continual NeRF methods like MEIL-NeRF [Chung et al.(2022)Chung, Lee, Baik, and Lee] and CLNeRF [Shin et al.(2017)Shin, Lee, Kim, and Kim]. As shown in Table 1 (left), on the Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun], C-NGP achieves the highest PSNR and SSIM comparable to SCARF ($0.02$ units lower) while avoiding additional parameters per scene. On Nerf Synthetic $360^{\circ}$ [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] (see Table 1 (right)), C-NGP shows competitive performance with SCARF. However, over scenes from the LLFF dataset [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar], C-NGP does not achieve the best scores due to the inherent struggle of Instant-NGP over the unbounded scenes. To support the claim, more empirical results are detailed in the supplementary material.
Qualitative. We provide a qualitative comparison of C-NGP with MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su], IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser], and vanilla NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] in Figure 2 and 3, over NeRF synthetic and Real LLFF datasets, respectively. Given its multi-scene modeling capability, we expect C-NGP’s performance to at least closely match per-scene optimized vanilla NeRF. Our results show that C-NGP achieves the best rendering quality on the NeRF Synthetic dataset and closely matches vanilla NeRF (which performs better than Instant-NGP) on the Real Forward-Facing dataset [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar]. Additionally, we evaluate the visual fidelity of different scenes rendered by C-NGP across the NeRF Synthetic [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng], Real Forward-Facing [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar], and Tanks and Temples [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun] in the supplementary.
Refer to caption Refer to caption
Figure 4: (Left) The figure illustrates how the proposed C-NGP method stores multiple edit styles within a single network. Each scene is trained separately, represented by $\theta^{i}$, and stores different edit styles for respective scenes. (Right) Effect of the difference in camera parameters across scenes in the offline setting.
Method Current Scene [L] Current Scene [L $\xrightarrow{}$ C] Current Scene [L $\xrightarrow{}$ C $\xrightarrow{}$ S] Current Scene [L $\xrightarrow{}$ C $\xrightarrow{}$ S $\xrightarrow{}$ D]
Lego Chair Ship Drums Lego Chair Ship Drums Lego Chair Ship Drums Lego Chair Ship Drums
Offline Sampling
35.62 x x x 34.59 34.63 x x 34.19 34.06 28.53 x 14.41 18.12 28.06 25.02
Online Sampling
35.62 x x x 34.81 34.82 x x 34.24 34.35 28.68 x 33.39 33.89 28.29 25.20
Table 2: Analysing progressive degradation in the rendering quality with offline and online sampling of scenes from the NeRF Synthetic $360^{\circ}$ dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]. The cell highlighted in red color shows high degradation in the PSNR value under offline sampling.
4.4 Additional Results and Analysis
Style Editing. To show the application of the proposed C-NGP framework, we adapt it for style editing in 3D scenes [Nguyen-Phuoc et al.(2022)Nguyen-Phuoc, Liu, and Xiao, Huang et al.(2021)Huang, Tseng, Saini, Singh, and Yang, Kopanas et al.(2021)Kopanas, Philip, Leimkühler, and Drettakis]. The current methods are only able to do single style transfer to any NeRF model at a time [Haque et al.(2023)Haque, Tancik, Efros, Holynski, and Kanazawa, Fujiwara et al.(2024)Fujiwara, Mukuta, and Harada], and for each new style, they have to retrain the NeRF model. The incremental nature of C-NGP also allows for accommodating multiple different styles within a scene. For each scene, we train a separate C-NGP network $\theta^{1}$, $\theta^{2}$, $\theta^{3}$, $\theta^{4}$ and for each $\theta^{i}$, we first learn the same scene for $K$ times incrementally using C-NGP and to transfer the style, we only fine-tune the MLP layer responsible for color (RGB). As shown in Figure 4 (left), we learn two different styles for each scene, i.e., each $\theta^{i}$ stores three different variations of the same scene, where the first is the original scene itself, and the next two are the edited scene styles which are vintage and anime. To generate consistent multi-view style scenes for training, we follow the work by Fujiwara et al. [Fujiwara et al.(2024)Fujiwara, Mukuta, and Harada], which uses fully fused attention and depth-conditioned ControlNet [Zhang et al.(2023b)Zhang, Rao, and Agrawala].
Single Dataset Mixed Dataset
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Groups 1 (8 Scenes) 35.889 0.978 29.40 0.94
Groups 2 (12 Scenes) 33.989 0.971 28.295 0.915
Groups 3 (16 Scenes) 32.448 0.964 27.190 0.906
Groups 4 (22 Scenes) 30.472 0.954 25.744 0.896
Experiments PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
XYZ | C + $SE(\theta,\phi)$ 24.108 0.812 0.201
XYZ + $\psi(C)$ + $SE(\theta,\phi)$ 12.966 0.813 0.460
XYZ | C + $\psi(C)$ + $SE(\theta,\phi)$ 24.492 0.824 0.184
XYZ | C + $\psi(C)$ + $SE(\theta,\phi)$ + $\psi(C)$ 25.927 0.880 0.139
Table 3: (Left) Analysing upper bound on the number of scenes C-NGP can accommodate with reasonable loss in rendering quality across single and mixed datasets. Single: $22$ scenes from BlenderNeRF dataset. Mixed: 14 scenes from BlenderNeRF + $8$ scenes from NeRF Synthetic $360^{\circ}$ dataset. (Right) Ablation on the configuration choice for conditioning Instant-NGP network [Müller et al.(2022)Müller, Evans, Schied, and Keller] - the backbone of C-NGP.
Sampling Strategy. To study the effect of how well C-NGP performs when in a continual paradigm with and without using the generative replay method (using stored images of previous scenes). Here, we define online sampling and offline sampling for the case with and without generative replay, respectively. As shown in Table 1 (bottom-right), the performance with online and offline sampling is nearly the same over the BlenderNeRF dataset, in fact, better than with offline sampling over the NeRF Synthetic dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]. We attribute such performance trends to the nature of camera parameters. Figure 4 (right) shows such misalignments due to differences in camera parameters that eventually cause a reduction in PSNR, leading to reduced performance in the offline sampling. To further emphasize this, Table 2 exclusively evaluates the rendering quality of scenes from the NeRF Synthetic dataset with these sampling methods.
Upper bound. Table 3 (left) attempts to demonstrate the number of scenes that can be accommodated within a given set of parameters with an acceptable loss in the rendering quality of the previously learned scenes. While we do not claim to avoid forgetting completely, most importantly, we succeed in slowing it down to the extent that we can reasonably model around $\sim 20$ different scenes in a single neural radiance field.
Design choices. We evaluated different strategies for concatenating pseudo labels and their sinusoidal positional encodings with scene coordinates and viewing directions as summarized in Table 3 (right). In our work, we use the configuration with the highest PSNR. More architectural design choices are detailed in the supplementary.
Supplementary. Due to space constraints, we elaborate: (a) quantitative comparison with other methods, (b) time analysis, (c) shared feature space visualization, and (d) additional discussion in the supplementary material. We urge the readers to kindly refer to the supplementary material as well for a holistic understanding of this work.
5 Conclusion
We introduced C-NGP, a continual learning framework for incremental multi-scene modeling within a single neural radiance field. C-NGP accommodates new scenes while reasonably preserving the previously learned information via scene conditioning and generative replay. Extensive empirical studies on synthetic and real datasets demonstrate that C-NGP achieves strong performance in multi-scene representation learning and outperforms existing methods over training and inference speed, offering real-time rendering capabilities. While at this stage, C-NGP incurs an acceptable loss in rendering quality while modeling multiple scenes, its performance in highly diverse scene settings and attempt towards zero forgetting warrants further exploration.
Appendix A Quantitative Comparison with Other Methods
The baseline methods report performance under two experimental settings: with and without finetuning. (a) Without finetuning: The model is trained on a set of scenes and evaluated on test scenes without modifying the learned parameters. (b) With finetuning: The model is first trained on a set of scenes, then fine-tuned on training images of a specific scene for a few epochs before evaluation on its unseen views. Although finetuning may seem redundant, as the model has already seen those scenes and will naturally perform better, we introduce it for a fair comparison similar to CP-NeRF [He et al.(2023)He, Liang, Xiao, Chen, and Chen], which also reports performance after finetuning on specific scenes.
To ensure a fair comparison with C-NGP, we determine the best setting for each baseline. Since MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] and IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] are not trained on scenes from NeRF Synthetic [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] or Real forward-facing [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar], we compare MVSNeRF and IBRNet with finetuning over these datasets. In contrast, NeRF*, CP-NeRF [He et al.(2023)He, Liang, Xiao, Chen, and Chen], and C-NGP are optimized on scenes from these datasets and are evaluated without fine-tuning.
We quantitatively compare C-NGP with the baselines under the appropriate settings on the NeRF Synthetic $360^{\circ}$ [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] and Real forward-facing (LLFF) [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar] datasets, as shown in Table 4 (left) and Table 4 (right), respectively. On the NeRF Synthetic $360^{\circ}$ dataset, C-NGP achieves the best SSIM and LPIPS scores under both settings. In terms of PSNR, it falls short by only $0.14$ units in the no-finetuning setting despite learning continually without any additional network for conditioning. This minor trade-off comes with the significant advantage of modeling multiple scenes incrementally without expanding the model size.
On the Real forward-facing dataset, as shown in Table 4 (right), C-NGP does not achieve the highest performance. This is primarily due to the limitations of the Instant-NGP backbone, which itself struggles with forward-facing scenes due to unbounded scenes, as observed when compared to vanilla NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] (see Table 4 (right), Row 4 & 5). However, upon finetuning, C-NGP surpasses Instant-NGP (see Table 4 (right), Row 8). It is important to note that Instant-NGP was chosen as the backbone to leverage its faster training benefits, and we also evaluated vanilla NeRF in the proposed continual setting to demonstrate its capability to model multiple scenes.
NeRF Synthetic 360
Method Finetuning PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeRF* x 21.75 0.84 0.16
IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] 28.14 0.94 0.07
MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] 27.07 0.93 0.17
CP-NeRF [He et al.(2023)He, Liang, Xiao, Chen, and Chen] x 29.54 0.92 0.09
C-NGP (ours) x 29.40 0.94 0.09
CP-NeRF [He et al.(2023)He, Liang, Xiao, Chen, and Chen] 31.77 0.95 0.06
C-NGP (ours) 32.08 0.95 0.06
Real Forward-Facing
Method Finetuning PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] 26.73 0.85 0.18
MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] 25.45 0.88 0.19
CP-NeRF [He et al.(2023)He, Liang, Xiao, Chen, and Chen] x 25.41 0.77 0.20
NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] per-scene opt. 26.50 0.81 0.25
Instant-NGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] per-scene opt. 24.98 0.78 0.24
C-NGP (ours) x 22.12 0.66 0.37
CP-NeRF [He et al.(2023)He, Liang, Xiao, Chen, and Chen] 27.23 0.81 0.14
C-NGP (ours) 25.19 0.76 0.25
Table 4: Quantitative comparison on NeRF Synthetic $360^{\circ}$ dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] (left) and Real Forward-Facing dataset [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar] (right). Left: Our continual framework shows competitive performance with multi-scene modeling and outperforms previous state-of-the-art methods. NeRF* indicates vanilla NeRF trained under the proposed continual setup. Right: We compare C-NGP with Instant-NGP (highlighted in red) and vanilla NeRF. C-NGP’s performance is constrained by Instant-NGP but improves with fine-tuning.
Method Per Scene [S, F, H, L] Fine-tune [S $\xrightarrow{}$ F] Fine-tune [S $\xrightarrow{}$ F $\xrightarrow{}$ H] Fine-tune [S $\xrightarrow{}$ F $\xrightarrow{}$ H $\xrightarrow{}$ L]
Ship Ficus Hotdog Lego Ship Ficus Hotdog Lego Ship Ficus Hotdog Lego Ship Ficus Hotdog Lego Avg.
MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] 21.27 19.60 22.44 18.90 12.61 17.72 x x 15.52 16.83 18.50 x 14.63 16.75 15.80 14.43 15.40
IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] 28.70 29.19 37.14 28.23 22.63 25.23 x x 23.98 24.63 35.16 x 23.97 24.71 33.49 30.72 28.22
C-NGP (ours)
(no fine-tuning)
30.14 33.95 37.38 35.62 28.86 32.03 x x 28.56 31.03 35.31 x 28.06 30.61 34.62 32.31 31.4
Table 5: Comparing the extent of information forgetting in MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su], IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser], and C-NGP across scenes in the NeRF Synthetic $360^{\circ}$ dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng].
Method
Training Time/
Complete Dataset
Fine-tuning Time/
Scene
Rendering Time/
Frame
NeRF* $\sim$12 days x $\sim$30 sec
MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] x $\sim$1 hour $\sim$14 sec
IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] x $\sim$9 hours $\sim$18 sec
CP-NeRF [He et al.(2023)He, Liang, Xiao, Chen, and Chen] >2 days x x
C-NGP (ours) $\sim$8 hours $\sim$10 mins $\sim$1.2 sec
Table 6: Training, fine-tuning, and rendering time comparison of C-NGP with other methods on the dataset NeRF Synthetic $360^{\circ}$. All time comparison experiments are done on NVIDIA RTX Quadro 5000 GPU. NeRF* indicates vanilla NeRF trained under the proposed continual learning setup.
Notably, if the training data of previously observed scenes is available, fine-tuning C-NGP on those scenes can further improve rendering quality. However, we see, even without finetuning, C-NGP delivers significantly better results than the baselines.
Forgetting Effect. Interestingly, Table 5 compares the extent of forgetting in MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] and IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] with C-NeRF (not fine-tuned) evaluated over the scenes in the NeRF Synthetic $360^{\circ}$ [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] dataset. The experimental setup for the forgetting experiment is that we first fine-tune a scene on MVSNeRF and IBRNet. Later, for each new scene, we use the previous scene’s fine-tuned weights and fine-tune them. For the proposed method, we follow the continual learning strategy only. No separate fine-tuning is done for C-NGP.
Appendix B Time Analysis
Training time and rendering speed are crucial factors when modeling multiple scenes. As shown in Table 6, NeRF requires $12$ days to train and $30$ seconds per frame to render, making it impractical. Other methods like MVSNeRF [Chen et al.(2021)Chen, Xu, Zhao, Zhang, Xiang, Yu, and Su] and IBRNet [Wang et al.(2021b)Wang, Wang, Genova, Srinivasan, Zhou, Barron, Martin-Brualla, Snavely, and Funkhouser] reduce training time but still have long fine-tuning and rendering times. In contrast, C-NGP trains in just $8$ hours, fine-tuning takes $10$ minutes per scene, and renders at $1.2$ seconds per frame, making it highly efficient for real-time applications.
Refer to caption Refer to caption
Figure 5: (Left) We combine different scenes of the NeRF Synthetic $360^{\circ}$ [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng] dataset and train Instant-NGP. The results show that neural hash encoding of scene coordinates may not be sufficient to preserve information across multiple scenes. (Right) The figure illustrates the representation space of C-NGP, visualized through a t-SNE map over scenes from NeRF Synthetic $360^{\circ}$ dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng].
Appendix C Shared Representation Space
As shown in Figure 5 (left), training Instant-NGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] on a mixed dataset without conditioning led to the intermixing of scene information. This confirms that neural hashing with only scene coordinates is insufficient for distinguishing scenes in a multi-scene setting. To achieve unique scene representations, additional conditioning is required.
Figure 5 (right) shows the t-SNE map [Van der Maaten and Hinton(2008)] of the learned multi-scene representation space of C-NGP across eight scenes from the NeRF Synthetic $360^{\circ}$ dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]. The overlapping scene clusters suggest that the parameter space is shared across multiple scenes, meaning a subset of parameters encodes shared information rather than being exclusive to a single scene.
To generate this visualization, we extract neural hash encodings from one image per scene, all captured from the same viewpoint. Since each image is $800\times 800$ pixels and many pixels correspond to the background, we divide it into $50\times 50$ patches, yielding 16 patches per image. Instead of extracting features from every pixel, we obtain MLP features from these patches across all eight scenes and visualize them using t-SNE [Van der Maaten and Hinton(2008)].
Require: $I_{new}$: list of new scene images
Require: $l_{new}$: pseudo label for new scene
Require: $L_{prev}$: list of pseudo labels for previous scenes
Require: $F_{\theta}$: NeRF network
Require: $C_{I_{new}}$: camera parameters of new scene
Require: $k$: number of images to render for previous scenes
/* Render previous scenes and store in $I_{prev}$/
Initialize $I_{prev}\xleftarrow{}\{\}$
for $l^{i}_{prev}\in L_{prev}$ do
       $render\_image\xleftarrow{}F_{\theta}(l^{i}_{prev},C_{I_{new}},k)$
       $I_{prev}$.append($render\_image$)
$T_{images}\xleftarrow{}concat(I_{new},I_{prev})$
  // train images list
$T_{labels}\xleftarrow{}concat(L_{new},L_{prev})$
  // train labels list
/* $TrainNeRF()$ train the $F_{\theta}$ on given scene list and labels */
$F_{\theta}^{updated}\xleftarrow{}TrainNeRF(F_{\theta},T_{images},T_{labels},C_{I_{new}})$
Output: Updated $F_{\theta}^{updated}$ on new scene
Algorithm 1 Learning new scenes using Conditional-cum-Continual NeRF (C-NGP)
Appendix D Architecture Choice
We conducted an ablation study to determine the most effective way to integrate multi-scene conditioning into the Instant-NGP [Müller et al.(2022)Müller, Evans, Schied, and Keller] architecture, and the results are summarized in the main paper Table $3$ (right).
The baseline, XYZ | C, feeds the pseudo label alongside $(x,y,z)$ coordinates into the neural hashing module. While straightforward, this approach relies solely on raw labels for scene differentiation. In XYZ + $\psi(C)$, we replaced the raw label with its positional encoding $\psi(C)$, but this led to lower performance, indicating that direct pseudo-label input is crucial for effective scene discrimination.
Adding both the raw label and its positional encoding (XYZ | C + $\psi(C)$) improved results, suggesting that combining these representations enhances scene modeling. The best performance was achieved with XYZ | C + $\psi(C)$ + $SE(\theta,\phi)$ + $\psi(C)$, where we further incorporated the spherical encoding of the viewing direction $SE(\theta,\phi)$ along with an additional instance of $\psi(C)$. This setup yielded the highest PSNR and SSIM scores and the lowest LPIPS, highlighting the benefit of jointly leveraging scene coordinates, viewing directions, and both raw and encoded labels.
Overall, our study shows that enriching multi-scene representations with positional and spherical encodings significantly improves scene modeling and retention.
Appendix E Additional Discussion
Empirically, we explored the upper bound on the number of scenes that a fixed number of parameters can model. Although information forgetting can be minimized further with an increase in the number of parameters and scenes, our goal was to maximally utilize the representation capacity of a single neural radiance field parameter. Interestingly, we show that online sampling under generative replay produces better results compared to offline sampling (algorithm 1). While the proposed framework can be viewed as taking the first steps toward multi-scene modeling, several open research avenues exist around it. An interesting future work could involve studying if the model learns useful scene priors during a continual learning process to aid in either faster learning over new scenes (i.e., convergence in fewer iterations or higher initial PSNR while learning to model a new scene) or can model new scenes with fewer number of images (few-shot learning). Moreover, dissecting C-NGP or the Instant-NGP backbone to analyze the kind of scene attributes handled by the model at each layer to increase physical interpretability is also an interesting avenue to explore. With the current exploration of the representative capacity of NeRFs and the aforementioned future directions, we believe that this work could serve as a primer for a new perspective on designing multi-scene continual neural radiance fields to continue accommodating new scenes without or with minimal loss of information of previously learned scenes.
Figure 6: Qualitative demonstration of the quality of rendered images across different views through C-NGP over the scenes from the NeRF Synthetic $360^{\circ}$ dataset [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng]. It is best viewed in PDF with Zoom.
Figure 7: Qualitative demonstration of the quality of rendered images across different views through C-NGP over the scenes from the Real Forward-Facing dataset [Mildenhall et al.(2019)Mildenhall, Srinivasan, Ortiz-Cayon, Kalantari, Ramamoorthi, Ng, and Kar]. It is best viewed in PDF with Zoom.
Figure 8: Qualitative demonstration of the quality of rendered images across different views through C-NGP over the scenes from the Tanks and Temples dataset [Knapitsch et al.(2017)Knapitsch, Park, Zhou, and Koltun].